SOS optics prepared to be hanged are moved from the South Flow Bench to S15 Clean Cabinet.
SRMU 03 (1-25-2010) specification summery E080460-05-D, older vintage SRM 01 and PRM 02 (need more specification)
There was one NOT MARKED SOS with two broken magnets on its face. This is labeled ???
This was done to prepare clean space for TIP-TILT drive- test set up.
The existing cable from 1X5 can reach only to the south end: from whitening filter to satelite amp. This will be good for future test of suspensions.
We need to make new cable from 1 X 1 to the south end = 40m long |